Derive the expression of the Q = ΔE of a nuclear reaction, hence, determine the Q value and the nature of the reaction F19 (H1,α) O16( 18.998405u, 15.994915u, 4.002603u, 1.007825u are masses of F19, O16, α, H1) respectively.



a ii). Discuss the primary processes by which gamma rays interact with matter



b). The Q value of the Na23 (n,α) F20 reaction is -5.4MeV. Determine the threshold energy of the neutrons for this reaction (22.9898u, 18.9964u, 4.0039u, 1.008665u are the masses of Na23, F20 α and n respectively).

1
Expert's answer
2022-02-03T09:03:14-0500

"^{19}F+^1H\to ^{16}O+^4He,"

"Q=\Delta E=(18.998405+1.007825-15.994915-4.002603)c^2=4.7~MeV,"


"^{23}Na+^1n\to^{20}F+^4He,"

"Q=\Delta E=(22.9898+1.008665-19.9964-4.0039)c^2=-5.3~MeV."